OpenStudy (anonymous):

. A solution of salt and water contains 60 grams of water per 120 milliliters of the solution. If 1 mole of water weighs 16 grams, _____ moles of water would be present in 40 milliliters of the solution

OpenStudy (zpupster):

how many moles in first solution??

OpenStudy (anonymous):

It doesnt say

OpenStudy (anonymous):

I wanna know how to solve these type of problems

OpenStudy (zpupster):

well 60/16 is the moles and 40/120 is a third, i would say the answer of 60/16 mult by a third.

OpenStudy (anonymous):

I still kinda dont get it

OpenStudy (zpupster):

these are ratio problems kind of like this 1mole of water/16 g= x mole of water/60 g = 60/16 = moles of water then that answer/120 ml = how many moles/40 = 40(that answer)/120= how many moles in 40

OpenStudy (anonymous):

But it doesnt say how much water each mole has. It says it weighs 16 grams

OpenStudy (zpupster):

yes there are 16 g/mol in water or 16 grams in 1 mole of water so how many moles in 60 hence 60/16 that is how many moles there are in 60 grams of water

OpenStudy (anonymous):

16 x 3.75 = 60 so 3.75 moles?

OpenStudy (zpupster):

ok now you know there are 3.75 moles per 120ml take that answer and plug it in the 2nd formula

OpenStudy (anonymous):

so do i do 3.75 / 40?

OpenStudy (zpupster):

per means divided by so 3.75/120 = x/40 or 40(3.75)/120

OpenStudy (anonymous):

it equals 0.0007812

OpenStudy (zpupster):

no punch in 40 times 3.75 and then divide by 120 you should get 1.25

OpenStudy (anonymous):

Ok. I used your first formula 3.75/120 = x/40. Now its asking me to "Round off your answer to two decimal places"

OpenStudy (zpupster):

it is already

OpenStudy (zpupster):

good job!! next time post this under chemistry.

OpenStudy (anonymous):

Its in my math class.

OpenStudy (anonymous):

But thank you so much!!

OpenStudy (anonymous):

so whats the answer